OpenStudy (anonymous):

Given the equation and solution, find the missing coefficient. 12x + 6 =____ x − 12 Solution: x = -9

jimthompson5910 (jim_thompson5910):

Let c be that missing coefficient, so 12x + 6 = c*x - 12 Plug in x = -9 to get 12(-9) + 6 = c*(-9) - 12 -108 + 6 = -9c - 12 -102 = -9c - 12 Now solve for c to find the missing coefficient.

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

good you got it